OpenStudy (anonymous):

what is 1/1+cot^2 theta expressed as a single trigonometric function?

OpenStudy (nincompoop):

1+cot^2 = ?

OpenStudy (anonymous):

my answer sheet says sin^2. i just need to know how to get there

OpenStudy (anonymous):

@nincompoop

OpenStudy (nincompoop):

well you need to know what the equivalent of 1+cot^2

OpenStudy (anonymous):

and that's sin^2. how did you get to the answer? I missed alot of the material, as I had an emergency, so I'm really confused

OpenStudy (nincompoop):

is 1+cot θ = sin^2?

OpenStudy (nincompoop):

I don't think so

OpenStudy (anonymous):

oh sorry, it's 1-cos^2

OpenStudy (nincompoop):

1+ cot^2 θ = ccs^2 θ and then 1/csc is the same as sin therefore, 1/csc^2 θ = sin^2 θ

OpenStudy (nincompoop):

review the pythagorean identities http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

1/(1 + (cos^2 x)/(sin^2 x) = (sin^2 x)/(sin^2 x + cos^2 x) = sin^2 x